# MFAI: A Scalable Bayesian Matrix Factorization Approach to Leveraging Auxiliary Information

### v2

@@ Appendix A Fitting the Single-Factor MFAI Model
where $\mathbf{z}\in\mathbb{R}^{N\times 1}$ is the latent loading, $\mathbf{w}\in\mathbb{R}^{M\times 1}$ is the latent factor, $\boldsymbol{\epsilon}\in\mathbb{R}^{N\times M}$ is a matrix of residual error terms, $F:\mathbb{R}^{C}\rightarrow\mathbb{R}$ is the unknown function, $F\left(\mathbf{X}\right)\in\mathbb{R}^{N\times 1}$ denotes $\left(F\left(\mathbf{X}_{1\cdot}\right),\dots,F\left(\mathbf{X}_{N\cdot}\right)\right)^{\mathrm{T}}$, $\mathbf{X}_{n\cdot}=\left(\mathbf{X}_{n1},\dots,\mathbf{X}_{nC}\right)^{\mathrm{T}}\in\mathbb{R}^{C\times 1}$ is the $n$-th row of $\mathbf{X}$ containing auxiliary information for the $n$-th row entity for $n=1,\dots,N$, and $\beta$ and $\tau$ are two precision parameters. Then, the joint probabilistic model becomes
@@
$\left(\widehat{\boldsymbol{\Theta}},\widehat{F}\left(\cdot\right)\right)$ $=\underset{\boldsymbol{\Theta},F\left(\cdot\right)}{\operatorname{arg}\,\operatorname{max}}\;\log{\rm Pr}\left(\mathbf{Y}\mid\boldsymbol{\Theta},F\left(\cdot\right)\right)$ (40)
@@ A.1 Approximate Bayesian Inference
[…]
@@ A.1.1 E-step
[…]
@@
[…]
@@ A.1.2 M-step
Here, we adopt the idea of the gradient boosting machine in each iteration to construct $F\left(\cdot\right)$. Suppose in the $(t-1)$-th step, the current estimate for $F\left(\cdot\right)$ is denoted as $F^{(t-1)}\left(\cdot\right)$. Specifically, boosting finds a minimizer of the loss function $\mathcal{L}\left(\boldsymbol{\mu},F\left(\cdot\right)\right)$ in a stage-wise manner by sequentially adding an update $f^{(t)}\left(\cdot\right)$ to $F^{(t-1)}\left(\cdot\right)$ in the $t$-th step
@@
$-\mathbf{H}^{-1}\left(\boldsymbol{\mu},F^{(t-1)}\left(\cdot\right)\right)\mathbf{G}\left(\boldsymbol{\mu},F^{(t-1)}\left(\cdot\right)\right)=\boldsymbol{\mu}-F^{(t-1)}\left(\mathbf{X}\right),$ (63)
@@
where $10<s<$ 1 is the so-called shrinkage parameter or learning rate.
@@ A.2 Missing Data in the Main Matrix
$\boldsymbol{\tau}_{nm}=\begin{cases}\tau,&\text{if $\left(n,m\right)\in\Omega^{\text{obs}}$},\\ 0,&\text{otherwise}.\end{cases}$ (69)
@@ Appendix B Fitting the Multi-Factor MFAI Model
1 set the maximum value of the number of factors $K_{\max}$ ;
@@
5       $\boldsymbol{\mu}_{k},a^{2}_{k};\boldsymbol{\nu}_{k},b^{2}_{k};\tau_{k},\beta_{k};F_{k}\left(\cdot\right)\leftarrow$ MFAI_SF$(\mathbf{Y},\mathbf{X})$ ;
@@ 11       $K\leftarrow K+1$ ;
14return $\widehat{\mathbf{Z}}=\left(\boldsymbol{\mu}_{1},\dots,\boldsymbol{\mu}_{K}\right),\widehat{\mathbf{W}}=\left(\boldsymbol{\nu}_{1},\dots,\boldsymbol{\nu}_{K}\right);a_{1}^{2},\dots,a_{K}^{2},b_{1}^{2},\dots,b_{K}^{2}$ ;
@@ C.1 Spatial and Temporal Information
1 Embryonic $4\ \mathrm{PCW}\leq\mathrm{Age}<8\ \mathrm{PCW}$
2 Early fetal $8\ \mathrm{PCW}\leq\mathrm{Age}<10\ \mathrm{PCW}$
3 Early fetal $10\ \mathrm{PCW}\leq\mathrm{Age}<13\ \mathrm{PCW}$
4 Early mid-fetal $13\ \mathrm{PCW}\leq\mathrm{Age}<16\ \mathrm{PCW}$
5 Early mid-fetal $16\ \mathrm{PCW}\leq\mathrm{Age}<19\ \mathrm{PCW}$
6 Late mid-fetal $19\ \mathrm{PCW}\leq\mathrm{Age}<24\ \mathrm{PCW}$
7 Late fetal $24\ \mathrm{PCW}\leq\mathrm{Age}<38\ \mathrm{PCW}$
8 Neonatal and early infancy $0\ \mathrm{M}\ \mathrm{(birth)}\leq\mathrm{Age}<6\ \mathrm{M}$
@@ 9 Late infancy $6\mathrm{M}\ \leq\mathrm{Age}<12\ \mathrm{M}$
10 Early childhood $1\ \mathrm{Y}\leq\mathrm{Age}<6\ \mathrm{Y}$
11 Middle and late childhood $6\ \mathrm{Y}\leq\mathrm{Age}<12\ \mathrm{Y}$
12 Adolescence $12\ \mathrm{Y}\leq\mathrm{Age}<20\ \mathrm{Y}$
13 Young adulthood $20\ \mathrm{Y}\leq\mathrm{Age}<40\ \mathrm{Y}$
14 Middle adulthood $40\ \mathrm{Y}\leq\mathrm{Age}<60\ \mathrm{Y}$
15 Late adulthood $60\ \mathrm{Y}\leq\mathrm{Age}$
@@ C.2 Differential Stability
Differential stability (DS) is defined as the tendency for a gene to exhibit reproducible differential expression relationships across brain structures (Shaw et al., 2011). DS can be measured by a variety of different means, such as the average Kendall-Tau correlation, average Spearman correlation, and average Euclidean difference (see Hawrylycz et al. (2015) Supplementary Analysis for comparable metrics). Here we use the average Pearson correlation as an example. Suppose the whole brain gene expression dataset contains $B$ individual brains, and we need to calculate the average Pearson correlation over $B(B-1)/2$ pairs of $B$ brains to measure DS between brains. We can represent the previous gene expression data matrix to a tensor $\mathbf{X}\in\mathbb{R}^{B\times S\times G}$, where $B$ denotes the individual brains, $S$ denotes the anatomic structures (refer to the neocortex areas in our analysis), and $G$ denotes the genes as in the main text. For each gene $g$ expressed in a pair of brains (indexed by $b_{1}$ and $b_{2}$) $\mathbf{X}_{b_{1}\cdot g}$ and $\mathbf{X}_{b_{2}\cdot g}$, let $\left\{\left(\mathbf{X}_{b_{1}1g},\mathbf{X}_{b_{2}1g}\right),\dots,\left(\mathbf{X}_{b_{1}Sg},\mathbf{X}_{b_{2}Sg}\right)\right\}$ be $S$ pairs expression level measurements in $S$ common anatomic structures and we computed the Pearson correlation $\rho_{S}\left(\mathbf{X}_{b_{1}\cdot g},\mathbf{X}_{b_{2}\cdot g}\right)$. Then the DS of gene $g$ is defined as the average $\rho_{S}$ over $B(B-1)/2$ pairs of brains. For the differentially expressed gene $g$, intuitively, $\mathbf{X}_{b_{1}s_{1}g}>\mathbf{X}_{b_{1}s_{2}g}$ is expected if $\mathbf{X}_{b_{2}s_{1}g}>\mathbf{X}_{b_{2}s_{2}g}$ or $\mathbf{X}_{b_{1}s_{1}g}<\mathbf{X}_{b_{1}s_{2}g}$ is expected if $\mathbf{X}_{b_{2}s_{1}g}<\mathbf{X}_{b_{2}s_{2}g}$, for any $b_{1},b_{2}\in\left\{1,\dots,B\right\}$ and any $s_{1},s_{2}\in\left\{1,\dots,S\right\}$. In summary, DS is a correlation-based metric that can be applied to assess the reproducibility of gene expression patterns across various structures in various individual brains and help reveal mesoscale genetic organization. Recent studies have shown that the genes with the highest DS are highly biologically relevant, with enrichment for brain-related annotations, disease associations, drug targets, and literature citations.

### v3

@@ Appendix A Fitting the Single-Factor MFAI Model
where $\mathbf{z}\in\mathbb{R}^{N\times 1}$ is the latent factor, $\mathbf{w}\in\mathbb{R}^{M\times 1}$ is the latent loading, $\boldsymbol{\epsilon}\in\mathbb{R}^{N\times M}$ is a matrix of residual error terms, $F:\mathbb{R}^{C}\rightarrow\mathbb{R}$ is the unknown function, $F\left(\mathbf{X}\right)\in\mathbb{R}^{N\times 1}$ denotes $\left(F\left(\mathbf{X}_{1\cdot}\right),\dots,F\left(\mathbf{X}_{N\cdot}\right)\right)^{\mathrm{T}}$, $\mathbf{X}_{n\cdot}=\left(\mathbf{X}_{n1},\dots,\mathbf{X}_{nC}\right)^{\mathrm{T}}\in\mathbb{R}^{C\times 1}$ is the $n$-th row of $\mathbf{X}$ containing auxiliary information for the $n$-th row entity for $n=1,\dots,N$, and $\beta$ and $\tau$ are two precision parameters. Then, the joint probabilistic model becomes
@@
$=\underset{\boldsymbol{\Theta},F\left(\cdot\right)}{\operatorname{arg}\,\operatorname{max}}\;\log\int{\rm Pr}\left(\mathbf{Y},\mathbf{z},\mathbf{w}\mid\boldsymbol{\Theta},F\left(\cdot\right)\right)\,d\mathbf{z}\,d\mathbf{w}.$
@@ A.1 Approximate Bayesian Inference
$\log{\rm Pr}\left(\mathbf{Y}\mid\boldsymbol{\Theta},F\left(\cdot\right)\right)$ $=\log\int{\rm Pr}\left(\mathbf{Y},\mathbf{z},\mathbf{w}\mid\boldsymbol{\Theta},F\left(\cdot\right)\right)\,d\mathbf{z}\,d\mathbf{w}$ (42)
$=\log\int q\left(\mathbf{z},\mathbf{w}\right)\frac{{\rm Pr}\left(\mathbf{Y},\mathbf{z},\mathbf{w}\mid\boldsymbol{\Theta},F\left(\cdot\right)\right)}{q\left(\mathbf{z},\mathbf{w}\right)}\,d\mathbf{z}\,d\mathbf{w}$
@@ A.1.1 E-step
$=\int q\left(\mathbf{z}\right)\frac{\log\widetilde{{\rm Pr}}\left(\mathbf{Y},\mathbf{z}\right)}{\log q\left(\mathbf{z}\right)}\,d\mathbf{z}+\text{const},$
@@
$=\int q\left(\mathbf{w}\right)\frac{\log\widetilde{{\rm Pr}}\left(\mathbf{Y},\mathbf{w}\right)}{\log q\left(\mathbf{w}\right)}\,d\mathbf{w}+\text{const},$
@@ A.1.2 M-step
Here, we adopt the idea of the gradient boosting machine in each iteration to construct $F\left(\cdot\right)$. Suppose in the (t1)(t-1)-th step, the current estimate for $F\left(\cdot\right)$ is denoted as $F^{(t-1)}\left(\cdot\right)$. Specifically, boosting finds a minimizer of the loss function $\mathcal{L}\left(\boldsymbol{\mu},F\left(\cdot\right)\right)$ in a stage-wise manner by sequentially adding an update $f^{(t)}\left(\cdot\right)$ to $F^{(t-1)}\left(\cdot\right)$ in the $t$-th step
@@
$\mathbf{H}^{-1}\left(\boldsymbol{\mu},F^{(t-1)}\left(\cdot\right)\right)\mathbf{G}\left(\boldsymbol{\mu},F^{(t-1)}\left(\cdot\right)\right)=\boldsymbol{\mu}-F^{(t-1)}\left(\mathbf{X}\right),$ (63)
@@
where $0<s<1$ is the so-called shrinkage parameter or learning rate.
@@ A.2 Missing Data in the Main Matrix
[…]
@@ Appendix B Fitting the Multi-Factor MFAI Model
1
21exset the maximum value of the number of factors $K_{\max}$ ;
@@
[…]
@@ 12 $K\leftarrow K+1$ ;
15return $\widehat{\mathbf{Z}}=\left(\boldsymbol{\mu}_{1},\dots,\boldsymbol{\mu}_{K}\right),\widehat{\mathbf{W}}=\left(\boldsymbol{\nu}_{1},\dots,\boldsymbol{\nu}_{K}\right);a_{1}^{2},\dots,a_{K}^{2},b_{1}^{2},\dots,b_{K}^{2}$ ;
@@ C.1 Spatial and Temporal Information
[…]
@@ 9 Late infancy $6\mathrm{M}\ \leq\mathrm{Age}<12\ \mathrm{M}$
[…]
@@ C.2 Differential Stability
Differential stability (DS) is defined as the tendency for a gene to exhibit reproducible differential expression relationships across brain structures (Shaw et al., 2011). DS can be measured by a variety of different means, such as the average Kendall-Tau correlation, average Spearman correlation, and average Euclidean difference (see Hawrylycz et al. (2015) Supplementary Analysis for comparable metrics). Here we use the average Pearson correlation as an example. Suppose the whole brain gene expression dataset contains $B$ individual brains, and we need to calculate the average Pearson correlation over B(B1)/2B(B-1)/2 pairs of $B$ brains to measure DS between brains. We can represent the previous gene expression data matrix to a tensor $\mathbf{X}\in\mathbb{R}^{B\times S\times G}$, where $B$ denotes the individual brains, $S$ denotes the anatomic structures (refer to the neocortex areas in our analysis), and $G$ denotes the genes as in the main text. For each gene $g$ expressed in a pair of brains (indexed by $b_{1}$ and $b_{2}$) $\mathbf{X}_{b_{1}\cdot g}$ and $\mathbf{X}_{b_{2}\cdot g}$, let $\left\{\left(\mathbf{X}_{b_{1}1g},\mathbf{X}_{b_{2}1g}\right),\dots,\left(\mathbf{X}_{b_{1}Sg},\mathbf{X}_{b_{2}Sg}\right)\right\}$ be $S$ pairs expression level measurements in $S$ common anatomic structures and we computed the Pearson correlation $\rho_{S}\left(\mathbf{X}_{b_{1}\cdot g},\mathbf{X}_{b_{2}\cdot g}\right)$. Then the DS of gene $g$ is defined as the average $\rho_{S}$ over B(B1)/2B(B-1)/2 pairs of brains. For the differentially expressed gene $g$, intuitively, $\mathbf{X}_{b_{1}s_{1}g}>\mathbf{X}_{b_{1}s_{2}g}$ is expected if $\mathbf{X}_{b_{2}s_{1}g}>\mathbf{X}_{b_{2}s_{2}g}$ or $\mathbf{X}_{b_{1}s_{1}g}<\mathbf{X}_{b_{1}s_{2}g}$ is expected if $\mathbf{X}_{b_{2}s_{1}g}<\mathbf{X}_{b_{2}s_{2}g}$, for any $b_{1},b_{2}\in\left\{1,\dots,B\right\}$ and any $s_{1},s_{2}\in\left\{1,\dots,S\right\}$. In summary, DS is a correlation-based metric that can be applied to assess the reproducibility of gene expression patterns across various structures in various individual brains and help reveal mesoscale genetic organization. Recent studies have shown that the genes with the highest DS are highly biologically relevant, with enrichment for brain-related annotations, disease associations, drug targets, and literature citations.